A map of the world is 7 feet wide and 3 feet tall. What is its perimeter?


1
Expert's answer
2021-08-26T13:25:17-0400

"l=7 \\

w=3 \\

P = 2 \times l + 2 \times w \\

= 2 \times 7 + 2 \times 3 \\

= 14 + 6 = 20 \;feet"